OpenStudy (anonymous):

calculus question find the points of discontinuity, and determine whether the discontinuties are removable f(x)=x^2-4/x^3-8

OpenStudy (anonymous):

factor factor factor then cancel

OpenStudy (anonymous):

how do you factor the bottom

OpenStudy (anonymous):

do you know how to factor \[x^3-8\]?

OpenStudy (anonymous):

you are going to need this one in calc, so best to memorize it \[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

in your case \(a=x,b=2\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

once you cancel, you have "removed" the discontinuity, so it will be "removable"

OpenStudy (anonymous):

so the removable disconuity would be 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

denominator is 0 at \(x=2\) so it is the function is discontinuous there but since if \(x\neq 2\) this whole thing is \[\frac{x+2}{x^2+2x+4}\] and since the denominator is never zero, you have only one discontinuity at \(x=2\) and you have just removed it

OpenStudy (anonymous):

yw